LASER EXCITATION SPECTROSCOPY OF SrSH AND CaSH

\underline{M. J. DICK}, Department of Physics, University of Waterloo, 200 University Ave. W., Waterloo, ON, N2L 3G1 Canada; P. M. SHERIDAN, J. G. WANG and P. F. BERNATH, Department of Chemistry, University of Waterloo, 200 University Ave. W. , Waterloo, ON, N2L 3G1 Canada.

The Ã^{2}A′ − X̃^{2}A′, B̃^{2}A′′ − X̃^{2}A′ and C̃^{2}A′ − X̃^{2}A′ transitions of SrSH along with the C̃^{2}A′ − X̃^{2}A′ transition of CaSH have been observed in our laser ablation spectrometer. These species are created via the reaction of strontium or calcium atoms with H_{2}S gas entrained in argon. The spectra are then measured at low and high resolution using laser excitation spectroscopy. For each of the observed electronic transitions, at least one K_{a} sub-band has been observed along with the 1 − 0 vibrational bands of the M–S stretch (M=Ca or Sr). The results of a preliminary fit of the high resolution data will be presented along with a comparison of the derived spectroscopic parameters for CaSH and SrSH.